# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 6. The unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$
Finite classical groups have at most one irreducible unipotent cuspidal representation. Recall that an irreducible representation of a finite reductive group $G^{F}$ is said to be unipotent if it occurs in $R_{T,1}$, where $T$ is an $F$-stable torus in $G$. If an irreducible unipotent cuspidal representation exists, then it is invariant under automorphisms of the underlying algebraic group. Additionally, we have
@@
So, the case of unipotent cuspidal representations for finite classical groups is a natural testing ground for Treumann–Venkatesh conjectures. In this section we treat the group ${\rm Sp}_{4}(\mathbb{F}_{q})$ which has a unique cuspidal unipotent representation for all $q$. Recall that ${\rm Sp}_{4}(\mathbb{F}_{q})$ is defined as the group of all $g\in{\rm GL}_{4}(\mathbb{F}_{q})$ such that $gJg^{t}=J$, where
@@
Let $F$ be the Frobenius endomorphism of ${\rm Sp}_{4}/\mathbb{F}_{q}$. (Note that $g^{t}$ is the transpose of $g$.) In the famous article [Sri68], B.Srinivasan discovered an irreducible cuspidal representation of the group ${\rm Sp}_{4}(\mathbb{F}_{q})$, which does not have Whittaker model and this representation is denoted by $\theta_{10}$. This representation $\theta_{10}$ is an irreducible unipotent cuspidal representation.
@@ 6.1. Mirabolic restriction of $\theta_{10}$
where $w=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $I_{2}$ is the $2\times 2$ identity matrix. Fix a non-trivial additive character $\psi:\mathbb{F}_{q}\rightarrow\overline{\mathbb{Q}}_{l}^{\times}$, and an element $\tau\in\mathbb{F}_{q}^{\times}\setminus(\mathbb{F}_{q}^{\times})^{2}$. We set
@@
Consider the character $\psi_{z}:U(\mathbb{F}_{q})\rightarrow\overline{\mathbb{Q}}_{l}^{\times}$, sending
@@
Then $\psi_{z}$ induces a character $\widetilde{\psi_{z}}:Y(\mathbb{F}_{q})\rightarrow\overline{\mathbb{Q}}_{\ell}^{\times}$, defined as
@@
for $s\in S(\mathbb{F}_{q})$ and $u\in U(\mathbb{F}_{q})$. Then, the induced representation ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ is irreducible, and, from [Gol78, Proposition 1], we get that the restriction ${\rm res}_{P(\mathbb{F}_{q})}(\theta_{10})$ is isomorphic to ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$. Moreover, the $-\ell$ reduction of ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ is irreducible, and this implies that the $-\ell$ reduction $r_{\ell}(\theta_{10})$ is also irreducible representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$.
6.2.
@@
Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. We denote by $\Pi$ and $\pi$ the irreducible cuspidal unipotent $\ell$-adic representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$ and ${\rm Sp}_{4}(\mathbb{F}_{q})$, respectively. By uniqueness of $\Pi$, we have $\Pi\simeq\Pi^{\sigma}$. Let $\eta$ be the additive character $\mathbb{F}_{q^{\ell}}$ given by composition $\psi\circ{\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$, where ${\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$ is the trace function. Since $\ell$ is odd, we have $\tau\in\mathbb{F}_{q^{\ell}}^{\times}\setminus(\mathbb{F}_{q^{\ell}}^{\times})^{2}$, and let $\widetilde{\eta}_{z}$ be the corresponding additive character of $U(\mathbb{F}_{q^{\ell}})$. Then, following the above constructions, we get an irreducible representation ${\rm Ind}_{Y(\mathbb{F}_{q^{l}})}^{P(\mathbb{F}_{q^{l}})}(\widetilde{\eta_{z}})$ of $P(\mathbb{F}_{q^{\ell}})$, and the restriction ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to ${\rm Ind}_{Y(\mathbb{F}_{q^{l}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$. We use the notations $\rho$ and $\rho_{\ell}$ for the induced representations ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ and ${\rm Ind}_{Y(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$, respectively.
6.2.1.
@@
Let $X$ and $X_{\ell}$ be the right coset spaces $Y(\mathbb{F}_{q})\backslash P(\mathbb{F}_{q})$ and $Y(\mathbb{F}_{q^{\ell}})\backslash P(\mathbb{F}_{q^{\ell}})$ respectively. Since $Y(\mathbb{F}_{q^{\ell}})$ is $\langle\sigma\rangle$ stable subgroup of $P(\mathbb{F}_{q^{\ell}})$, we have the long exact sequence of non-abelian cohomology
$0\rightarrow Y(\mathbb{F}_{q})\rightarrow P(\mathbb{F}_{q})\rightarrow X_{\ell}^{\sigma}\rightarrow H^{1}(\sigma,Y(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\sigma,P(\mathbb{F}_{q^{\ell}}))$ (6.1)
@@
Note that $S(\mathbb{F}_{q^{\ell}})$ is a $\langle\sigma\rangle$ stable subgroup of $Y(\mathbb{F}_{q^{\ell}})$ with $S(\mathbb{F}_{q^{\ell}})^{\sigma}=S(\mathbb{F}_{q})$. Then it follows from the long exact sequence of non-abelian cohomology
$0\rightarrow S(\mathbb{F}_{q})\rightarrow Y(\mathbb{F}_{q})\rightarrow U(\mathbb{F}_{q})\rightarrow H^{1}(\sigma,S(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\sigma,Y(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\sigma,U(\mathbb{F}_{q^{\ell}}))$
that the cohomology $H^{1}(\sigma,S(\mathbb{F}_{q^{\ell}}))$ and $H^{1}(\sigma,Y(\mathbb{F}_{q^{\ell}}))$ are equal. Let $S$ be the algebraic group such that $S^{F}=S(\mathbb{F}_{q})$, where $F$ is the Frobenius endomorphism of ${\rm Sp}_{4}/\mathbb{F}_{q}$ induced by $x\mapsto x^{q}$. The group $S$ is precisely given by
@@
where $g^{\ast}$ denotes the transposition of $g$ with respect to the second diagonal. Note that the determinants of the elements in $S$ are ±$\pm 1$, and the connected component of $S$, denoted by $S_{1}$, consists of matrices of determinant $1$. Then we have the following long exact sequence
$0\rightarrow H^{0}(\sigma,S_{1})\rightarrow H^{0}(\sigma,S)\rightarrow H^{0}(\sigma,\mathbb{Z}/2\mathbb{Z})\rightarrow H^{1}(\sigma,S_{1})\rightarrow H^{1}(\sigma,S)\rightarrow H^{1}(\sigma,\mathbb{Z}/2\mathbb{Z}),$
Using [DM91, Proposition 4.2.11], we get that $H^{1}(\sigma,S_{1})$ is trivial. Also, $H^{1}(\sigma,\mathbb{Z}/2\mathbb{Z})$ is trivial since $\ell$ is odd and $\sigma$ acts trivially on $\mathbb{Z}/2\mathbb{Z}$. Then it follows from the above exact sequence that $H^{1}(\sigma,S)$ is trivial, and so is $H^{1}(\sigma,Y(\mathbb{F}_{q^{\ell}}))$. Thus, the exact sequence (6.1) gives the equality $X_{\ell}^{\sigma}=X$.
Theorem 6.1.
@@
for all $x\in P(\mathbb{F}_{q^{\ell}})$ and $f\in\rho_{\ell}$. Let $\mathcal{L}$ be the $P(\mathbb{F}_{q^{\ell}})\rtimes\langle\sigma\rangle$ stable lattice in $\rho_{\ell}$ consisting of all $\overline{\mathbb{Z}}_{\ell}$-valued functions in $\rho_{\ell}$. Since $\widetilde{\eta_{z}}(\sigma(y))=\widetilde{\eta_{z}}(y)$ for $y\in Y(\mathbb{F}_{q})$, the Tate cohomology group $\widehat{H}^{1}(\widetilde{\eta}_{z})$ is trivial. Then it follows from isomorphism (2.1) that
@@
Since the $-\ell$ reduction $r_{\ell}(\rho_{\ell})$ is irreducible, any $P(\mathbb{F}_{q^{\ell}})$ stable $\Lambda$-lattice is homothetic to $\mathcal{L}$. Thus, we get that $\widehat{H}^{1}(\mathcal{L})$ is trivial for any $P(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ stable lattice $\mathcal{L}$ in $\rho_{\ell}$. ∎
6.2.2.
The next lemma will show that the zeroth Tate cohomology of the $-\ell$ reduction of $\Pi$ is irreducible and cuspidal.
Lemma 6.2.
Let $\Pi$ be the irreducible unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$. Then the zeroth Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is irreducible and cuspidal.
@@
First note that $r_{\ell}(\Pi)$ is irreducible. Further, it follows from the exactness of the Jacquet functor and [Vig96, Chapter I, Section 9.3] that $r_{\ell}(\Pi)$ is cuspidal. We now show that the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is cuspidal. Let $P$ be a proper parabolic subgroup of ${\rm Sp}_{4}$ with unipotent radical $U$. Using Proposition 3.2, we get an injection
$r_{{\rm Sp}_{4}(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widehat{H}^{0}(r_{\ell}(\Pi)))\hookrightarrow\widehat{H}^{0}(r_{{\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(r_{\ell}(\Pi))).$
Since $r_{\ell}(\Pi)$ is cuspidal, the Jacquet module $r_{{\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(r_{\ell}(\Pi))$ is trivial, and hence it follows from the above injection that $r_{{\rm Sp}_{4}(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widehat{H}^{0}(r_{\ell}(\Pi)))$ is trivial. Thus, $\widehat{H}^{0}(r_{\ell}(\Pi))$ is cuspidal.
It follows from (2.1) that the restriction to $P(\mathbb{F}_{q})$ map
@@
$f\longmapsto{\rm res}_{P(\mathbb{F}_{q})}(f)$
induces an isomorphism of $P(\mathbb{F}_{q})$ representations
@@
Since ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\rho_{\ell})$, the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$, being isomorphic to $r_{\ell}(\rho)^{(\ell)}$, is an irreducible representation of $P(\mathbb{F}_{q})$, and hence of ${\rm Sp}_{4}(\mathbb{F}_{q})$. This shows the lemma. ∎
6.3.
Next, we prove that the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to the Frobenius twist of the $-\ell$ reduction $r_{\ell}(\pi)$. Let $Q$ be the Klingen parabolic subgroup of ${\rm Sp}_{4}$ consisting of matrices $(g_{ij})\in{\rm GL}_{4}$ subject to the conditions $g_{21}=g_{31}=g_{42}=g_{43}=0$. Let $L$ be the Levi subgroup of $Q$ and $H$ be the unipotent radical of $Q$. Then $L(\mathbb{F}_{q})$ is isomorphic to $\mathbb{F}_{q}^{\times}\times{\rm SL}_{2}(\mathbb{F}_{q})$, and $H(\mathbb{F}_{q})$ consists of matrices of the form

### v2

@@ 7. The unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$
Finite classical groups can have at most one irreducible unipotent cuspidal representation. Let $G$ be a connected reductive group defined over $\mathbb{F}_{q}$, and let $F:G\rightarrow G$ be the Frobenius endomorphism. Recall that an irreducible representation of a finite reductive group $G^{F}$ is said to be unipotent if it occurs in a Deligne–Lusztig virtual character ±$\pm R_{T,1}$, where $T$ is an $F$-stable torus in $G$. When $G$ is a classical group, if an irreducible unipotent cuspidal representation exists, then it is invariant under automorphisms of the underlying algebraic group. Additionally, we have
@@
So, the case of unipotent cuspidal representations for finite classical groups is a natural testing ground for Treumann–Venkatesh conjectures. In this section, we treat the group ${\rm Sp}_{4}(\mathbb{F}_{q})$ which has a unique cuspidal unipotent representation for all $q$. Recall that ${\rm Sp}_{4}(\mathbb{F}_{q})$ is defined as the group of all $g\in{\rm GL}_{4}(\mathbb{F}_{q})$ such that $gJg^{t}=J$ (Note that $g^{t}$ is the transpose of $g$), where
@@
Let $F$ be the Frobenius endomorphism of ${\rm Sp}_{4}/\mathbb{F}_{q}$. In the famous article [Sri68], B. Srinivasan discovered an irreducible cuspidal representation of the group ${\rm Sp}_{4}(\mathbb{F}_{q})$, which does not have Whittaker model, and this representation is denoted by $\theta_{10}$. The representation $\theta_{10}$ is an unipotent cuspidal representation. In what follows we assume that $\ell\neq 2$.
@@ 7.1. Mirabolic restriction of $\theta_{10}$
where $w=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $I_{2}$ is the $\times 2$ 2 identity matrix. Fix a non-trivial additive character $\psi:\mathbb{F}_{q}\rightarrow\overline{\mathbb{Q}}_{\ell}^{\ast}$, and an element $\tau\in\mathbb{F}_{q}^{\times}\setminus(\mathbb{F}_{q}^{\times})^{2}$. We set
@@
Consider the character $\psi_{z}:U(\mathbb{F}_{q})\rightarrow\overline{\mathbb{Q}}_{\ell}^{\ast}$, sending
@@
Then $\psi_{z}$ induces a character $\widetilde{\psi_{z}}:Y(\mathbb{F}_{q})\rightarrow\overline{\mathbb{Q}}_{\ell}^{\ast}$, defined as
@@
for $s\in S(\mathbb{F}_{q})$ and $u\in U(\mathbb{F}_{q})$. The induced representation ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ is irreducible and from [Gol78, Proposition 1] we get that the restriction ${\rm res}_{P(\mathbb{F}_{q})}(\theta_{10})$ is isomorphic to ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$. Moreover, the $-\ell$ reduction of ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ is irreducible and hence the $-\ell$ reduction $r_{\ell}(\theta_{10})$ is also irreducible.
7.2.
@@
Let $\sigma$ be the Frobenius element in ${\Gamma=\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. We denote by $\Pi$ and $\pi$ the unipotent $\ell$-adic cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$ and ${\rm Sp}_{4}(\mathbb{F}_{q})$, respectively. By uniqueness of $\Pi$ we have $\Pi\simeq\Pi^{\sigma}$. Let $\eta$ be the additive character of $\mathbb{F}_{q^{\ell}}$ given by the composition $\psi\circ{\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$, where ${\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$ is the trace function. Since $\ell$ is odd, we have $\tau\in\mathbb{F}_{q^{\ell}}^{\times}\setminus(\mathbb{F}_{q^{\ell}}^{\times})^{2}$. Let $\eta_{z}$ and $\widetilde{\eta}_{z}$ be the corresponding characters of $U(\mathbb{F}_{q^{\ell}})$ and $Y(\mathbb{F}_{q^{\ell}})$, respectively. Then, following the above constructions, we get an irreducible $P(\mathbb{F}_{q^{\ell}})$ representation ${\rm Ind}_{Y(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$ such that the restriction ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to ${\rm Ind}_{Y(\mathbb{F}_{q^{l}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$. We use the notations $\rho$ and $\rho_{\ell}$ for the representations ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ and ${\rm Ind}_{Y(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$, respectively.
7.2.1.
@@
Let $X$ and $X_{\ell}$ be the coset spaces $Y(\mathbb{F}_{q})\backslash P(\mathbb{F}_{q})$ and $Y(\mathbb{F}_{q^{\ell}})\backslash P(\mathbb{F}_{q^{\ell}})$, respectively. Since $Y(\mathbb{F}_{q^{\ell}})$ is $\Gamma$ stable subgroup of $P(\mathbb{F}_{q^{\ell}})$, we have the long exact sequence of non-abelian cohomology ([Ser, Appendix, Proposition 1])
$0\rightarrow Y(\mathbb{F}_{q})\rightarrow P(\mathbb{F}_{q})\rightarrow X_{\ell}^{\sigma}\rightarrow H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\Gamma,P(\mathbb{F}_{q^{\ell}}))$ (7.1)
@@
Note that $S(\mathbb{F}_{q^{\ell}})$ is a $\Gamma$ stable subgroup of $Y(\mathbb{F}_{q^{\ell}})$ with $S(\mathbb{F}_{q^{\ell}})^{\sigma}=S(\mathbb{F}_{q})$, and we get the following the long exact sequence of non-abelian cohomology
$0\rightarrow S(\mathbb{F}_{q})\rightarrow Y(\mathbb{F}_{q})\rightarrow U(\mathbb{F}_{q})\rightarrow H^{1}(\Gamma,S(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\Gamma,U(\mathbb{F}_{q^{\ell}})).$
From the above exact sequence, we observe that $H^{1}(\Gamma,S(\mathbb{F}_{q^{\ell}}))$ is isomorphic to $H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))$. The group $S$ is precisely given by
@@
where $g^{\ast}$ is the transposition of $g$ with respect to the second diagonal. Note that the determinant map $S\rightarrow\{\pm 1\}$, given by $g\mapsto{\rm det}(g)$ is surjective. The kernel of this determinant map, denoted by $S_{1}$, is the connected component of $S$. Then we have the following long exact sequence
$0\rightarrow H^{0}(\Gamma,S_{1})\rightarrow H^{0}(\Gamma,S)\rightarrow H^{0}(\Gamma,\mathbb{Z}/2\mathbb{Z})\rightarrow H^{1}(\Gamma,S_{1})\rightarrow H^{1}(\Gamma,S)\rightarrow H^{1}(\Gamma,\mathbb{Z}/2\mathbb{Z}).$
Using [DM91, Proposition 4.2.11], we get that $H^{1}(\Gamma,S_{1})$ is trivial. Since $\ell$ is odd, we have $H^{1}(\Gamma,\mathbb{Z}/2\mathbb{Z})=\{0\}$. Then it follows from the above exact sequence that $H^{1}(\Gamma,S)$ is trivial, and so is $H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))$. Thus, the exact sequence (7.1) gives the equality $X_{\ell}^{\sigma}=X$.
Theorem 7.2.1.
@@
for all $x\in P(\mathbb{F}_{q^{\ell}})$ and $f\in\rho_{\ell}$. Let $\mathcal{L}$ be the $P(\mathbb{F}_{q^{\ell}})\rtimes\Gamma$ stable lattice in $\rho_{\ell}$ consisting of $\Lambda$-valued functions in $\rho_{\ell}$. Since $\widetilde{\eta_{z}}(\sigma(y))=\widetilde{\eta_{z}}(y)$ for $y\in Y(\mathbb{F}_{q})$, the Tate cohomology group $\widehat{H}^{1}(\widetilde{\eta}_{z})$ is trivial. Then it follows from the isomorphism (2.1) that
@@
Since the $-\ell$ reduction $r_{\ell}(\rho_{\ell})$ is irreducible, any $P(\mathbb{F}_{q^{\ell}})$ stable lattice is homothetic to $\mathcal{L}$. Thus, we get that $\widehat{H}^{1}(\mathcal{L})$ is trivial for any $\Lambda[P(\mathbb{F}_{q^{\ell}})\rtimes\Gamma]$ stable lattice $\mathcal{L}$ in $\rho_{\ell}$. ∎
7.2.2.
Next, we show that the zeroth Tate cohomology of the $-\ell$ reduction of $\Pi$ is irreducible and cuspidal.
Lemma 7.2.2.
Let $\Pi$ be the irreducible unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$. Then the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is irreducible and cuspidal.
@@
Note that $r_{\ell}(\Pi)$ is irreducible. Further, it follows from the exactness of the Jacquet functor and [Vig96, Chapter I, Section 9.3] that $r_{\ell}(\Pi)$ is also cuspidal. By Corollary 3.3.3, we get that $\widehat{H}^{0}(r_{\ell}(\Pi))$ is cuspidal.
To prove the irreducibility of $\widehat{H}^{0}(r_{\ell}(\Pi))$, we consider the following restriction to $P(\mathbb{F}_{q})$ map
@@
$f\longmapsto{\rm res}_{P(\mathbb{F}_{q})}(f).$
The above map factors through $\widehat{H}^{0}(r_{\ell}(\rho_{\ell}))$ and induces the following $P(\mathbb{F}_{q})$ isomorphism (see Subsection 2.2)
@@
Since the restriction ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\rho_{\ell})$, the Tate cohomology $\widehat{H}^{0}(r_{\ell}(\Pi))$, being $P(\mathbb{F}_{q})$ isomorphic to $r_{\ell}(\rho)^{(\ell)}$, is an irreducible representation of $P(\mathbb{F}_{q})$, and hence of ${\rm Sp}_{4}(\mathbb{F}_{q})$. This shows the lemma. ∎
7.3.
Next, we show that the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$. Let $Q$ be the Klingen parabolic subgroup of ${\rm Sp}_{4}$ consisting of matrices $(g_{ij})\in{\rm GL}_{4}$ with $g_{21}=g_{31}=g_{42}=g_{43}=0$. Let $L$ be the Levi subgroup of $Q$ and $H$ be the unipotent radical of $Q$. Then $L(\mathbb{F}_{q})$ is isomorphic to $\mathbb{F}_{q}^{\times}\times{\rm SL}_{2}(\mathbb{F}_{q})$, and $H(\mathbb{F}_{q})$ consists of matrices of the form